The sum of two numbers is 45. The larger number, y, is 6 less than twice the smaller number, x. What is the smaller number x? Also what is the equation?
Answer by lwsshak3(11628)

The sum of two numbers is 45. The larger number, y, is 6 less than twice the smaller number, x. What is the smaller number x? Also what is the equation?
**
let x=smaller number
y=2x-6=larger number
x+y=45
x+2x-6=45 (equation)
3x=39
x=13=smaller number